OpenStudy (aravindg):

A ball is thrown from the ground level so as to just clear a wall 4 m high at a distance of 4m and falls at a distance of 14 m from the wall.Find magnitude and direction of velocity of the ball??

OpenStudy (amistre64):



OpenStudy (amistre64):

-4.9 t^2 +A sin(Vi) t = 0 A cos(Vi) t = 28 would be equations to work with i think

OpenStudy (amistre64):

t = 28/A cos(Vi) -4.9 A sec(Vi)^2 +28 Atan(Vi) = 0 maybe? but i gotta get to class ..... good luck with it :)

OpenStudy (aravindg):

2.A particle is projected with a velocity v=ai+bj.Find the radius of curvature of trajectory of particle at the (1)point of projection (2)highest point

OpenStudy (anonymous):

\[v = ai + b j\] hmm Aravind a, b are constants right...but theres a problem the y velocity should depend on x or else when I differentiate it I get zero which implies no acceleration while in projectile motion the acceleration on y axis is -g...

OpenStudy (anonymous):

hmm if you want to know the Radius for general projectile, I can get that ...wait a minute let me post the solution

OpenStudy (aravindg):

i need the answer without using differentiation

OpenStudy (anonymous):

Let \(v_i\) be the speed with which projectile is projected. Then at Highest Point \(v_y=0\) And \(v_x\) Remains constant as there is no acceleration on x axis or horizontal component of velocity. \(\theta\) be the angle at which projectile is projected. \[v_icos\theta=v_{x}\] hmm Now....\(\frac{v^2}{R}=a\) \[R = \frac{v^2}{a}\] \[a=g\]\[v=v_x=v_i*cos\theta\]\[R=\frac{v_i^2*cos^2\theta}{g}\]

OpenStudy (anonymous):

The Answer I posted is for Radius at Highest Point ....

OpenStudy (aravindg):

k R= radius and not range ryt??

OpenStudy (anonymous):

Yea..R is Radius

OpenStudy (aravindg):

nikvist???

OpenStudy (aravindg):

...........

OpenStudy (aravindg):

........................................

OpenStudy (nikvist):

\[y=y(x)\quad;\quad A(0,0)\quad,\quad B(4,4)\quad,\quad C(18,0)\] \[y=px^2+qx+r\] \[A(0,0)\quad:\quad 0=r\] \[y=x(px+q)\quad\Rightarrow\quad 4=4(4p+q)\quad,\quad 0=18(18p+q)\quad\Rightarrow\quad q=-18p\] \[4p+q=1=4p-18p\quad\Rightarrow\quad p=-\frac{1}{14}\quad,\quad q=\frac{18}{14}\quad\Rightarrow\quad y=-\frac{1}{14}x^2+\frac{18}{14}x\] \[y'=-\frac{2}{14}x+\frac{18}{14}\quad,\quad\tan\theta=y'(0)=\frac{9}{7}\quad\Rightarrow\quad\theta=\arctan\frac{9}{7}=52.125^{o}\] \[d_\max=\frac{v^2_0}{g}\sin{2\theta}\quad,\quad v_0=\sqrt{\frac{gd_\max}{\sin{2\theta}}}=13.63\,\,m/s\quad(g\approx10m/s^2)\]

OpenStudy (nikvist):

math+physics method

OpenStudy (anonymous):

Cool Method .... Thanks I've never seen anyone doing this before .... : )..

OpenStudy (anonymous):

Method is Super COOL.......

OpenStudy (aravindg):

can u explain wat u did??

OpenStudy (nikvist):

I find equation of trajectory y(x)

OpenStudy (aravindg):

what is the answer of first question???i am not clear what amistre said

OpenStudy (nikvist):

answer is in my first reply

OpenStudy (aravindg):

k now 2nd question

OpenStudy (aravindg):

3.Two cars A and B start off to a race on a straight path with initial velocities 8m/s and 5 m/s respectively.Car A moves with uniform acceleration of 1 m/s^2 and car B moves with uniform acceleration 1.1m/s^2.If both the cars reach the winning post together,find length of the race track.Also find which of the two cars was ahead 10 s before the finish.

OpenStudy (nikvist):

\[\vec{v}=a\vec{i}+b\vec{j}\quad,\quad\tan\theta=\frac{b}{a}\] (1) point of projection \[R_1=\frac{v^2_1}{a_{1\perp}}=\frac{a^2+b^2}{g\cdot\cos\theta}=\frac{a^2+b^2}{g}\sqrt{1+\tan^2\theta}=\frac{a^2+b^2}{g}\sqrt{1+\frac{b^2}{a^2}}=\frac{\left(a^2+b^2\right)^{3/2}}{ga}\] (2) highest point \[R_1=\frac{v^2_2}{a_{2\perp}}=\frac{a^2}{g}\]

OpenStudy (aravindg):

ishaan got the answeer???

OpenStudy (anonymous):

umm No..I am on it ...

OpenStudy (aravindg):

nikvist???

OpenStudy (aravindg):

wow nikvist plz help in 3rd question

OpenStudy (aravindg):

4.A car accelerates from rest with a constant acceleration \[\alpha\] on a straight road.After gaining a velocity v,the car moves with that velocity for sometime.Then the car decelerates with a retardation \[\beta\].If total distance covered by car is equal to s ,find total time of its motion.

OpenStudy (nikvist):

\[s_A=v_{0A}t+\frac{1}{2}a_At^2\quad,\quad s_B=v_{0B}t+\frac{1}{2}a_Bt^2\] \[s_A=s_B\quad\Rightarrow\quad v_{0A}t+\frac{1}{2}a_At^2=v_{0B}t+\frac{1}{2}a_Bt^2\quad\Rightarrow\quad (v_{0A}-v_{0B})t=\frac{1}{2}(a_B-a_A)t^2\] \[t=2\frac{v_{0A}-v_{0B}}{a_B-a_A}=2\frac{3\,\,m/s}{0.1\,\,m/s^2}=60s\] \[s=8\cdot 60+\frac{1}{2}\cdot 3600=2280\,\,m\] A was ahead, because it started fast

OpenStudy (aravindg):

i became a big fan of u .Now 4th one

OpenStudy (aravindg):

5.A racing motor speeds up in a straight line in a lake ,from rest .Referring to the accleration -displacement graph for the speeding boat,find its speed when it passes a raft at a distance of s from starting point.

OpenStudy (nikvist):

\[s=s_1+s_2+s_3\] \[s_1=\frac{v^2}{2\alpha}\quad,\quad t_1=\frac{v}{\alpha}\] \[s_3=\frac{v^2}{2\beta}\quad,\quad t_3=\frac{v}{\beta}\] \[s_2=s-s_1-s_3=s-\frac{v^2}{2\alpha}-\frac{v^2}{2\beta}\quad\Rightarrow\quad t_2=\frac{s_2}{v}=\frac{s}{v}-\frac{v}{2\alpha}-\frac{v}{2\beta}\] \[t=t_1+t_2+t_3=\frac{v}{\alpha}+\frac{s}{v}-\frac{v}{2\alpha}-\frac{v}{2\beta}+\frac{v}{\beta}=\frac{s}{v}+\frac{v}{2\alpha}+\frac{v}{2\beta}\]

OpenStudy (aravindg):

you rock SUPER COOL now 5th one

OpenStudy (aravindg):

nikviiiiiiiiiiii

OpenStudy (nikvist):

\[s=\frac{v^2}{2\alpha}+\frac{v^2}{2\beta}=v^2\frac{\alpha+\beta}{2\alpha\beta}\] \[v=\sqrt{\frac{2\alpha\beta}{\alpha+\beta}\,s}\]

OpenStudy (aravindg):

6.Prove that path of a projectile with respect to another projectile is a straight line...

OpenStudy (aravindg):

thx now 6

OpenStudy (aravindg):

7.Two bodies are projected simultaneously with mutually perpendicular velocities each of magnitude v0=\[\sqrt{2*g*h}\] from top and bottom of a cliff of height h=8m.If the bodies collide in air,how far from the foot of the cliff does the collision occur?

OpenStudy (nikvist):

\[\vec{u}(t)=u_{0x}\vec{i}+(u_{0y}-gt)\vec{j}\] \[\vec{v}(t)=v_{0x}\vec{i}+(v_{0y}-gt)\vec{j}\] \[\vec{v}_{rel}=\vec{v}-\vec{u}=v_{0x}\vec{i}+(v_{0y}-gt)\vec{j}-u_{0x}\vec{i}-(u_{0y}-gt)\vec{j}=(v_{0x}-u_{0x})\vec{i}+(v_{0y}-u_{0y})\vec{j}\] \[\vec{v}_{rel}\neq\vec{v}_{rel}(t)\] one direction=straight line

OpenStudy (aravindg):

wow now 7th

OpenStudy (aravindg):

8.A body is thrown at an angle theta with horizontal such that it attains a speed equal to \[\sqrt{2/3}\] times the speed of projection when the body is at half of its maximum height.Find the angle theta.

OpenStudy (nikvist):

pause after 7th